if i traveled around the equator at 60 mph and 200 miles above the equator. How long would it take?

Answer by Alan3354(69443)

if i traveled around the equator at 60 mph and 200 miles above the equator. How long would it take?
----------------
r = radius of the Earth in miles
C = 2*pi*(r+200)
-------
d = r*t
t = d/r = 2*pi*(r+200)/60 hours
= (pi*r + 200pi)/30 hours